OpenStudy (anonymous):

write 21/48 in simplest form 21/48 7/48 7/16 21/16 and please explain!

myininaya (myininaya):

do 21 and 48 have any common factors? and do you know what a factors is?

OpenStudy (anonymous):

no i forgot:\

myininaya (myininaya):

Example: if i asked you to factor 9, one way to do is write 3*3. 3 is a factor of 9 just like you could write the factorization of 9 as 9*1 and you could say 9 is a factor and 1 is a factor

myininaya (myininaya):

we could write the prime factorization for both 21 and 48

OpenStudy (anonymous):

the common factor is 7?

myininaya (myininaya):

prime numbers are integers that can only be factored at 1 times that number like 2,3,5,7,11,13,17,19,23,29, and so on composite numbers are integers are not prime numbers like the ones in this fraction 21 and 48 are composite because we can write 21=3*7 48=2*2*2*2*3

myininaya (myininaya):

they only have a 3 in common

OpenStudy (anonymous):

i see

myininaya (myininaya):

so we have \[\frac{21}{48}=\frac{\not{3} \cdot 7}{16 \cdot \not{3}}\]

myininaya (myininaya):

see nothing but what they had in common canceled

myininaya (myininaya):

so 21/48 can be reduced to 7/16

OpenStudy (anonymous):

okay am getting it now! thanks(:

myininaya (myininaya):

np :)